# Reinforcement Fine-Tuning Naturally Mitigates Forgetting in Continual Post-Training

### v1

@@ Abstract
Continual post-training (CPT) is a popular and effective technique for adapting foundation models like multimodal large language models to specific and ever-evolving downstream tasks. While existing research has primarily concentrated on methods like data replay, model expansion, or parameter regularization, the fundamental role of the learning paradigm within CPT remains largely unexplored. This paper presents a comparative analysis of two core post-training paradigms: supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT), investigating their respective impacts on knowledge retention during CPT. Our experiments are conducted on a benchmark comprising seven diverse multimodal tasks, utilizing Qwen2.5-VL-7B-Instruct as the base model for continual post-training. The investigation yields two significant findings: (1) When continuously learning on downstream tasks, SFT leads to catastrophic forgetting of previously learned tasks. In contrast, RFT inherently preserves prior knowledge and achieve performance comparable to multi-task training. (2) RFT successfully protects and even enhances the model’s general knowledge on standard benchmarks (e.g., MMMU and MMLU-Pro). Conversely, SFT degrades general model capabilities severely. Further analysis shows that explicit mechanisms, such as KL penalty and chain-of-thought reasoning, are not the primary factors. Instead, we find that the implicit regularization inherent to RFT is a key factor in mitigating forgetting. Finally, we propose a rollout-based instance filtering algorithm to improve the stability and efficiency of RFT. Our comprehensive study demonstrates the superiority of RFT as a robust paradigm for continual post-training.
Figure 1: Comparison of performance retention between SFT and RFT in continual post-training. We plot the performance relative to the initial training performance on each task as learning progresses through a sequence of tasks. (a) SFT exhibits catastrophic forgetting; (b) RFT maintains performance on previously learned tasks.
@@ 1 Introduction
Recently, reinforcement fine-tuning (RFT), which optimizes models based on feedback from generated outputs, has significantly advanced foundation model post-training (Chu et al., 2025; Shao et al., 2024; Guo et al., 2025a). In this work, we conduct the first investigation into whether SFT or RFT is the more suitable paradigm for CPT, focusing on knowledge preservation for both specific downstream tasks and general capabilities. Experimentally, we continually fine-tune the Qwen2.5-VL-7B-Instruct model (Bai et al., 2025b) on a benchmark comprising diverse multimodal tasks covering various domains. To fully reflect the knowledge preservation ability, we evaluate forgetting on both learned specific tasks and general benchmarks such as MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a).
@@
To understand how RFT mitigates forgetting during CPT, we conduct additional experiments with the popular and representative group relative policy optimization (GRPO) framework (Shao et al., 2024). We analyze the impact of KL divergence penalty and chain-of-thought (CoT) reasoning (Wei et al., 2022) on forgetting mitigation. Particularly, the KL divergence penalty prevents the policy from changing too drastically, similar to the well-known knowledge distillation in continual learning (Li & Hoiem, 2017). However, we reveal that KL penalty and CoT mechanisms are not the primary factors, and instead, the implicit regularization introduced by RFT significantly contributes to forgetting mitigation. Concretely, the expected magnitude of the gradient update, when measured in the parameter subspace sensitive to old tasks, is inherently more conservative than that of SFT. This conservatism is naturally scaled by the variance of the reward signal, effectively creating a data-dependent regularization. Last but not least, we observe that the learning process of RFT can be highly inefficient. Thus, we introduce a rollout-based instance filtering algorithm that enhances the stability of GRPO while still being an excellent knowledge protector.
@@
    We present a comprehensive analysis of the forgetting mitigation effects of SFT and RFT during continual post-training of MLLMs, demonstrating that RFT naturally preserves not only the performance of learned downstream tasks but also general model capabilities.
@@ 3 Preliminaries
where $\beta>0$. ReMax (Li et al., 2023c) use the reward of a greedy decoding response $a_{\text{greedy}}$ as a baseline. For a single sampled response $a$, the objective is to maximize:
$\mathcal{J}_{\text{ReMax}}(\theta)=\mathbb{E}_{x,a\sim\pi_{\theta}}\left[(r(x,a)-r(x,a_{\text{greedy}}))\log\pi_{\theta}(a|x)\right].$ (4)
This adaptive baseline helps to normalize rewards and reduce gradient variance. To further reduce variance, RLOO (Ahmadian et al., 2024) generates $n$ samples $\{a_{1},\dots,a_{n}\}$ and uses the average reward of the other $n-1$ samples as a baseline for sample $a_{i}$:
@@ 4.1 Experimental Setup
All experiments employ full-parameter fine-tuning for both SFT and RFT to ensure comprehensive capability assessment. Experiments of SFT are conducted using the llamafactory (Zheng et al., 2024) framework, with a learning rate of $51e-$ 5 and a batch size of 24. RFT methods (GRPO, ReMax, and RLOO) are implemented using the easyR1 (Zheng et al., 2025) framework, building upon Verl (Sheng et al., 2024). A consistent configuration is applied across RFT methods to ensure an equitable comparison: a learning rate of $61e-$ 6, a rollout batch size of 512, a sampling temperature of 1.0, with KL-divergence coefficient $\beta=0.01$. Specifically, GRPO is implemented adhering to its foundational methodology, with a group size set to 8. ReMax followed its core algorithm, and RLOO adopted the official Hugging Face algorithm.
@@ 4.3 Finding 2: RFT Protects and Enhances General Capabilities
Our experiments reveal that SFT causes significant “base model degradation” (Liu et al., 2025a). SFT induces a severe performance drop of $\downarrow$16.9on the challenging MMLU-Pro benchmark (47.5% $\rightarrow$ 30.6%). Crucially, this is not merely an artifact of sequential learning; even multi-task SFT (MTL (SFT)), which trains on all data simultaneously, still causes a severe drop of $\downarrow$14.3on the same benchmark. A similar trend is evident on MMMU, where SFT and MTL (SFT) cause performance to decline by $\downarrow$12.0and $\downarrow$4.3respectively. This demonstrates that the SFT paradigm itself appears harmful to the model’s foundational capabilities.
@@
In contrast to the capability decay observed across all SFT methods, the RFT paradigm effectively safeguards the model’s general abilities. GRPO, in particular, often enhances these abilities. For instance, GRPO improves performance on MMMU by $\uparrow$2.1(52.1% $\rightarrow$ 54.2%). Crucially, RFT also improves model general capabilities, with GRPO improving the POPE score by $\uparrow$1.9(86.6% $\rightarrow$ 88.5%) and reducing the tendency for hallucination. This clear difference highlights that RFT is a more robust paradigm for continual post-training.
@@ 5.1 The Roles of CoT and KL Penalty
5.2 Theoretical Insight: Implicit Regularization from Reward Variance
To explore the difference between SFT and RFT, we analyze their gradient updates from the perspective of continual learning, revealing that RFT’s forgetting mitigation stems from an implicit regularization mechanism.
We formalize this by defining a measure of how much a gradient update interferes with knowledge from past tasks. Following foundational work in continual learning (Kirkpatrick et al., 2017), we use the Fisher Information Matrix (FIM) as a theoretical tool to quantify parameter sensitivity.
@@ 5.2 Theoretical Insight: Implicit Regularization from Reward Variance
Theorem 5.2 (RFT’s Implicit Regularization Effect).
@@
$\mathbb{E}_{a\sim\pi_{\theta_{k-1}}}[\mathcal{R}(g_{\text{RFT}}(a))]=\text{Var}_{a\sim\pi_{\theta_{k-1}}}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\mathcal{E},$ (10)
where $\mathcal{E}$ is an error term that is small when the policy is reasonably accurate. The term $\text{Var}[r(x_{k},a)]$ is bounded by $1/4$ for normalized rewards.
The full proof is provided in Appendix B. Theorem 5.2 reveals that the expected RFT risk is scaled by the variance of reward. When the model is uncertain about the correct answer for a given prompt, the reward variance is bounded, leading to an update in parameter subspaces important for old tasks, as shown in Figure 4.
@@ 5.3 RIF-RFT: Enhancing Stability and Efficiency of RFT
The mechanism is formalized in Algorithm 1. For each instance in a new task’s dataset $\mathcal{D}_{k}$, we perform a small number of policy rollouts. If at least one of these rollouts produces a response with a reward greater than a minimal threshold $\tau$, we classify the instance and retain it in $\mathcal{D}_{k}^{\text{filt}}$. As shown in Table 5, RIF-RFT achieves competitive performance (57.5% of AvgAcc) and strong knowledge retention (-4.5% of FM), which is superior to SFT with a FM of -10.4%. Meanwhile, RIF-RFT use only 48.4% of training data. This demonstrates a compelling trade-off between efficiency and robustness.
Input: New task training set: $\mathcal{D}_{k}=\{(x_{i},y_{i}^{*})\}_{i=1}^{M}$; Current model policy: $\pi_{\theta}$; Number of rollouts per input: $N$; Reward threshold: $\tau$
for each sample $(x_{i},y_{i}^{*})\in\mathcal{D}_{k}$ do
       Initialize $R_{\max}\leftarrow-\infty$;
       for $j=1$ to $N$ do
             Sample a response: $y_{ij}\sim\pi_{\theta}(\cdot\mid x_{i})$;
             Compute reward: $R(y_{ij})$;
             Update: $R_{\max}\leftarrow\max(R_{\max},R(y_{ij}))$;
      if $R_{\max}>\tau$ then
             Add $(x_{i},y_{i}^{*})$ to $\mathcal{D}_{k}^{\text{filt}}$;
Output: Updated model $\pi_{\theta^{\prime}}$
Algorithm 1 Rollout-based Instance Filtering for RFT (RIF-RFT)
@@ 6 Conclusion
This work presents a comprehensive investigation into the role of the fundamental learning paradigm in continual post-training for MLLMs. Our central finding is that RFT naturally mitigates the catastrophic forgetting that plagues the standard SFT. Through extensive experiments, we demonstrate that while SFT leads to severe degradation of both previously learned task-specific skills and general capabilities, RFT paradigms inherently preserve those knowledge, achieving performance comparable to an offline multi-task learning setting. Our in-depth analysis reveals that this superiority is not due to explicit mechanisms like CoT or KL regularization, but is instead a result of the implicit regularization from reward-variance-weighted updates in RFT. Finally, we introduce RIF-RFT, an efficient instance filtering method that improves the stability and sample efficiency of RFT without compromising its robustness. This research suggests that RFT is not merely an alternative but a fundamentally more suitable paradigm for the continual and lifelong adaptation of foundation models.
@@ Appendix A Dataset Information
Appendix B Proof and Technical Details for Theorem 5.2
We provide the detailed derivation for Theorem 5.2, which establishes the relationship between the forgetting risks of RFT and SFT.
Theorem 5.2.
@@
$\mathbb{E}_{a\sim\pi_{\theta_{k-1}}}[\mathcal{R}(g_{\text{RFT}}(a))]=\text{Var}_{a\sim\pi_{\theta_{k-1}}}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\mathcal{E},$ (11)
@@
$I(a)\triangleq(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a|x_{k})).$ (12)
@@
$\mathbb{E}_{a\sim\pi_{\theta}}[A(x_{k},a)^{2}]=\text{Var}_{a\sim\pi_{\theta}}[r(x_{k},a)].$ (13)
@@
$\mathbb{E}[A^{2}]$ $=\mathbb{E}[(r-b)^{2}]=\mathbb{E}[r^{2}]-2b\mathbb{E}[r]+b^{2}.$ (14)
Since $b=\mathbb{E}[r]$, this simplifies to $\mathbb{E}[r^{2}]-2(\mathbb{E}[r])^{2}+(\mathbb{E}[r])^{2}=\mathbb{E}[r^{2}]-(\mathbb{E}[r])^{2}=\text{Var}[r]$. ∎
Proof of Theorem 5.2.
Let us analyze the forgetting risks at parameters $\theta=\theta_{k-1}$ for a single data point $(x_{k},a_{k}^{*})$. The analysis proceeds in three steps.
Step 1: SFT Forgetting Risk.
@@
$\mathcal{R}(g_{\text{SFT}})=g_{\text{SFT}}^{\top}F_{k-1}g_{\text{SFT}}=(\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k}))=I(a_{k}^{*}).$ (15)
Step 2: Expected RFT Forgetting Risk.
@@
$\mathbb{E}[\mathcal{R}(g_{\text{RFT}})]$ $=\mathbb{E}_{a\sim\pi_{\theta}}\left[(g_{\text{RFT}}(a))^{\top}F_{k-1}(g_{\text{RFT}}(a))\right]$ (16)
$=\mathbb{E}_{a\sim\pi_{\theta}}\left[A(x_{k},a)^{2}\cdot(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))\right]$ (17)
$=\mathbb{E}_{a\sim\pi_{\theta}}\left[A(x_{k},a)^{2}\cdot I(a)\right].$ (18)
Step 3: Relating RFT and SFT Risks.
@@
$\mathbb{E}[\mathcal{R}(g_{\text{RFT}})]$ $=\mathbb{E}[A(a)^{2}]\cdot\mathbb{E}[I(a)]+\text{Cov}(A(a)^{2},I(a))$ (19)
@@
$=\text{Var}[r(x_{k},a)]\cdot I(a_{k}^{*})+\underbrace{\text{Var}[r(x_{k},a)]\delta+\epsilon_{1}}_{\text{Error Term }\mathcal{E}}$ (22)
$=\text{Var}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\mathcal{E}.$ (23)
This completes the proof. The error term $\mathcal{E}$ is small when both $\epsilon_{1}$ and $\delta$ are small, which holds when the policy is reasonably well-behaved and the advantage signal is not strongly correlated with parameter importance. The primary relationship demonstrates that the expected forgetting risk of RFT is modulated by the reward variance, providing an implicit regularization against catastrophic forgetting. ∎

### v2

@@ Abstract
Continual post-training (CPT) is a popular and effective technique for adapting foundation models like multimodal large language models to specific and ever-evolving downstream tasks. While existing research has primarily concentrated on methods like data replay, model expansion, or parameter regularization, the fundamental role of the learning paradigm within CPT remains largely unexplored. This paper presents a comparative analysis of two core post-training paradigms: supervised fine-tuning (SFT) and reinforcement fine-tuning (RFT), investigating their respective impacts on knowledge retention during CPT. Our experiments are conducted on a benchmark comprising seven diverse multimodal tasks, utilizing Qwen2.5-VL-7B-Instruct as the base model for continual post-training. The investigation yields two significant findings: (1) When continuously learning on downstream tasks, SFT leads to catastrophic forgetting of previously learned tasks. In contrast, RFT inherently preserves prior knowledge and achieve performance comparable to multi-task training. (2) RFT successfully protects and even enhances the model’s general knowledge on standard benchmarks (e.g., MMMU and MMLU-Pro). Conversely, SFT degrades general model capabilities severely. Further analysis reveals that this stability is not primarily due to explicit mechanisms like KL penalty or chain-of-thought reasoning. Instead, we identify an implicit regularization mechanism inherent to RFT as a key contributing factor. Our theoretical analysis suggests that RFT’s gradient updates are naturally scaled by the reward variance, acting as a data-dependent regularizer that inherently protects previously acquired knowledge. Finally, we propose a rollout-based instance filtering algorithm to enhance the stability and efficiency of RFT. Our comprehensive study demonstrates the superiority of RFT as a robust paradigm for continual post-training.
Figure 1: Comparison of performance retention between SFT and RFT in continual post-training. We plot the performance on each task, normalized relative to its initial post-training peak, as the model learns through a sequence of multimodal tasks. (a) SFT exhibits classic catastrophic forgetting, where performance on previously learned tasks degrades dramatically as new tasks are introduced. (b) By contrast, RFT demonstrates remarkable stability, maintaining high performance on prior tasks throughout the entire sequence. This suggests an inherent forgetting-mitigation property within the RFT paradigm. Further details on the experimental setup can be found in Section 4.
@@ 1 Introduction
Recently, reinforcement fine-tuning (RFT), which optimizes models based on feedback from generated outputs, has significantly advanced foundation model post-training (Chu et al., 2025; Shao et al., 2024; Guo et al., 2025a). To the best of our knowledge, this work presents the first direct comparative investigation into whether SFT or RFT is the more suitable paradigm for CPT, focusing on knowledge preservation for both specific downstream tasks and general capabilities. Experimentally, we continually fine-tune the Qwen2.5-VL-7B-Instruct model (Bai et al., 2025b) on a benchmark comprising diverse multimodal tasks covering various domains. To fully reflect the knowledge preservation ability, we evaluate forgetting on both learned specific tasks and general benchmarks such as MMMU (Yue et al., 2024), MMLU-Pro (Wang et al., 2024b), and POPE (Li et al., 2023a).
@@
To understand how RFT mitigates forgetting during CPT, we conduct additional experiments with the popular and representative group relative policy optimization (GRPO) framework (Shao et al., 2024). We analyze the impact of KL divergence penalty and chain-of-thought (CoT) reasoning (Wei et al., 2022) on forgetting mitigation. Particularly, the KL divergence penalty prevents the policy from changing too drastically, similar to the well-known knowledge distillation in continual learning (Li & Hoiem, 2017). However, our analysis indicates that these explicit mechanisms are not the primary drivers of forgetting mitigation. We instead attribute this phenomenon to an implicit regularization effect within RFT. We offer a theoretical perspective suggesting that RFT’s updates are inherently more conservative in parameter subspaces sensitive to prior tasks. This conservatism is naturally scaled by the variance of the reward signal, creating a data-dependent regularization that dampens updates on uncertain samples, thus protecting established knowledge. Last but not least, we observe that the learning process of RFT can be highly inefficient. Thus, we introduce a rollout-based instance filtering algorithm that enhances the stability of GRPO while still being an excellent knowledge protector.
@@
    We present the first comprehensive analysis of the forgetting mitigation effects of SFT and RFT during continual post-training of MLLMs, demonstrating that RFT naturally preserves not only the performance of learned downstream tasks but also general model capabilities.
@@ 3 Preliminaries
where $\beta>0$. ReMax (Li et al., 2023c) use the reward of a greedy decoding response $a$ as a baseline. For a single sampled response $a$, the objective is to maximize:
$\mathcal{J}_{\text{ReMax}}(\theta)=\mathbb{E}_{x,a\sim\pi_{\theta}}\left[(r(x,a)-r(x,a))\log\pi_{\theta}(a|x)\right].$ (4)
This adaptive baseline helps to normalize rewards and reduce gradient variance. To further reduce variance, RLOO (Ahmadian et al., 2024) generates $n$ samples $\{a_{1},\dots,a_{n}\}$ and uses the average reward of the other n1n-1 samples as a baseline for sample $a_{i}$:
@@ 4.1 Experimental Setup
All experiments employ full-parameter fine-tuning for both SFT and RFT to ensure comprehensive capability assessment. Experiments of SFT are conducted using the llamafactory (Zheng et al., 2024) framework, with a learning rate of 1e51e-5 and a batch size of 24. RFT methods (GRPO, ReMax, and RLOO) are implemented using the easyR1 (Zheng et al., 2025) framework, building upon Verl (Sheng et al., 2024). A consistent configuration is applied across RFT methods to ensure an equitable comparison: a learning rate of 1e61e-6, a rollout batch size of 512, a sampling temperature of 1.0, with KL-divergence coefficient $\beta=0.01$. Specifically, GRPO is implemented adhering to its foundational methodology, with a group size set to 8. ReMax followed its core algorithm, and RLOO adopted the official Hugging Face algorithm.
@@ 4.3 Finding 2: RFT Protects and Enhances General Capabilities
Our experiments reveal that SFT causes significant “base model degradation” (Liu et al., 2025a). SFT induces a severe performance drop of $\downarrow$16.9% on the challenging MMLU-Pro benchmark (47.5% $\rightarrow$ 30.6%). Crucially, this is not merely an artifact of sequential learning; even multi-task SFT (MTL (SFT)), which trains on all data simultaneously, still causes a severe drop of $\downarrow$14.3% on the same benchmark. A similar trend is evident on MMMU, where SFT and MTL (SFT) cause performance to decline by $\downarrow$12.0% and $\downarrow$4.3% respectively. This demonstrates that the SFT paradigm itself appears harmful to the model’s foundational capabilities.
@@
In contrast to the capability decay observed across all SFT methods, the RFT paradigm effectively safeguards the model’s general abilities. GRPO, in particular, often enhances these abilities. For instance, GRPO improves performance on MMMU by $\uparrow$2.1% (52.1% $\rightarrow$ 54.2%). Crucially, RFT also improves model general capabilities, with GRPO improving the POPE score by $\uparrow$1.9% (86.6% $\rightarrow$ 88.5%) and reducing the tendency for hallucination. This clear difference highlights that RFT is a more robust paradigm for continual post-training.
@@ 5.1 The Roles of CoT and KL Penalty
5.2 Implicit Regularization from Reward Variance
To build intuition for the empirical resilience of RFT, we analyze its gradient dynamics in the context of continual learning. Our analysis suggests that RFT’s forgetting mitigation stems from an implicit regularization mechanism, where the learning signal itself modulates the update strength. To explore this intuition, we adopt the concept of forgetting risk from continual learning theory (Kirkpatrick et al., 2017), using the Fisher Information Matrix (FIM) as a tool to quantify parameter sensitivity to past tasks. This allows us to conceptually link the structure of RFT’s gradients to knowledge retention.
@@ 5.2 Implicit Regularization from Reward Variance
The following proposition establishes a conceptual link between the expected forgetting risk of an RFT update and that of an SFT update, highlighting the central role of reward variance.
Proposition 5.2 (RFT’s Implicit Regularization Effect).
@@
$\mathbb{E}_{a\sim\pi_{\theta_{k-1}}}[\mathcal{R}(g_{\text{RFT}}(a))]\approx\text{Var}_{a\sim\pi_{\theta_{k-1}}}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}}),$ (10)
where the approximation holds when an error term $\mathcal{E}$, capturing second-order effects, is small. The term $\text{Var}[r(x_{k},a)]$ is bounded by $1/4$ for normalized rewards.
The full proof is provided in Appendix B. Proposition 5.2 offers an intuition: the expected impact of an RFT update on prior knowledge is not fixed but is dynamically scaled by the reward variance. For an uncertain sample where the model generates diverse responses with high reward variance, the update magnitude in sensitive directions is naturally dampened, thus protecting established knowledge. Conversely, for samples where the model produces consistently high-reward responses (low variance), the update is more aggressive. This inherent, data-dependent regularization mechanism contrasts sharply with SFT’s uniform, high-variance gradients, offering a compelling explanation for the stability observed in our experiments and illustrated in Figure 4.
@@ 5.3 RIF-RFT: Enhancing Stability and Efficiency of RFT
The mechanism is formalized in Algorithm 1 in Appendix C. For each instance in a new task’s dataset $\mathcal{D}_{k}$, we perform a small number of policy rollouts. If at least one of these rollouts produces a response with a reward greater than a minimal threshold $\tau$, we classify the instance and retain it in $\mathcal{D}_{k}^{\text{filt}}$. As shown in Table 5, RIF-RFT achieves competitive performance (57.5% of AvgAcc) and strong knowledge retention (-4.5% of FM), which is superior to SFT with a FM of -10.4%. Meanwhile, RIF-RFT use only 48.4% of training data. This demonstrates a compelling trade-off between efficiency and robustness.
@@ 6 Conclusion
This work presents a comprehensive investigation into the role of the fundamental learning paradigm in continual post-training for MLLMs. Our central finding is that RFT naturally mitigates the catastrophic forgetting that plagues the standard SFT. Through extensive experiments, we demonstrate that while SFT leads to severe degradation of both previously learned task-specific skills and general capabilities, RFT paradigms inherently preserve those knowledge, achieving performance comparable to an offline multi-task learning setting. Our analysis suggests this superiority stems not from explicit mechanisms like CoT or KL regularization, but from an implicit regularization effect inherent to RFT. We provide a theoretical perspective that attributes this stability to reward-variance-scaled updates, which naturally protect previously acquired knowledge by moderating learning on uncertain samples. Finally, we introduce RIF-RFT, an efficient instance filtering method that improves the stability and sample efficiency of RFT without compromising its robustness. This research suggests that RFT is not merely an alternative but a fundamentally more suitable paradigm for the continual and lifelong adaptation of foundation models.
Ethics Statement
This research focuses on the fundamental learning paradigms for continual post-training of Multimodal Large Language Models. All experiments were conducted on publicly available and well-established academic benchmarks. Our work did not involve human subjects, private data, or generation of personally identifiable information.
Reproducibility Statement
To ensure the reproducibility of our findings, we provide comprehensive details throughout the paper and in the appendix. The source code is available in the supplementary material. Our experiments are based on the publicly available Qwen2.5-VL-7B-Instruct model. All implementation details are documented in Section 4 and the Appendix. We use standard public datasets, with detailed descriptions provided in Section 4 and the Appendix.
@@ Appendix A Dataset Information
Appendix B Proof and Technical Details for Proposition 5.2
We provide the detailed derivation for Proposition 5.2, which establishes the relationship between the forgetting risks of RFT and SFT.
Proposition 5.2.
@@
$\mathbb{E}_{a\sim\pi_{\theta_{k-1}}}[\mathcal{R}(g_{\text{RFT}}(a))]=\text{Var}_{a\sim\pi_{\theta_{k-1}}}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\mathcal{E}$
@@
$I(a)\triangleq(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))$
@@
$\mathbb{E}_{a\sim\pi_{\theta}}[A(x_{k},a)^{2}]=\text{Var}_{a\sim\pi_{\theta}}[r(x_{k},a)]$
@@
$\mathbb{E}[A^{2}]$ $=\mathbb{E}[(r-b)^{2}]=\mathbb{E}[r^{2}]-2b\mathbb{E}[r]+b^{2}$
Since $b=\mathbb{E}[r]$, this simplifies to $\mathbb{E}[r^{2}]-2(\mathbb{E}[r])^{2}+(\mathbb{E}[r])^{2}=\mathbb{E}[r^{2}]-(\mathbb{E}[r])^{2}=\text{Var}[r]$
Proof of Proposition 5.2.
Let us analyze the forgetting risks at parameters $\theta=\theta_{k-1}$ for a single data point $(x_{k},a_{k}^{*})$.
@@
$\mathcal{R}(g_{\text{SFT}})=g_{\text{SFT}}^{\top}F_{k-1}g_{\text{SFT}}=(\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a_{k}^{*}|x_{k}))=I(a_{k}^{*})$
@@
$\mathbb{E}[\mathcal{R}(g_{\text{RFT}})]$ $=\mathbb{E}_{a\sim\pi_{\theta}}\left[(g_{\text{RFT}}(a))^{\top}F_{k-1}(g_{\text{RFT}}(a))\right]$
$=\mathbb{E}_{a\sim\pi_{\theta}}\left[A(x_{k},a)^{2}\cdot(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))^{\top}F_{k-1}(\nabla_{\theta}\log\pi_{\theta}(a|x_{k}))\right]$
$=\mathbb{E}_{a\sim\pi_{\theta}}\left[A(x_{k},a)^{2}\cdot I(a)\right]$ (11)
@@
$\mathbb{E}[\mathcal{R}(g_{\text{RFT}})]$ $=\mathbb{E}[A(a)^{2}]\cdot\mathbb{E}[I(a)]+\text{Cov}(A(a)^{2},I(a))$
@@
$=\text{Var}[r(x_{k},a)]\cdot I(a_{k}^{*})+\text{Var}[r(x_{k},a)]\delta+\epsilon_{1}$
$=\text{Var}[r(x_{k},a)]\cdot\mathcal{R}(g_{\text{SFT}})+\underbrace{\text{Var}[r(x_{k},a)]\delta+\epsilon_{1}}_{\mathcal{E}}$
This completes the proof. The error term $\mathcal{E}=\text{Var}[r(x_{k},a)]\delta+\epsilon_{1}$ is small under reasonable conditions. Specifically, $\delta$ is small when the current policy is not drastically different from one that produces the ground-truth response, a condition met after initial task adaptation. $\epsilon_{1}$ is small if there is no systematic correlation between a response’s quality (reflected in its advantage) and its gradient’s impact on prior tasks, which is a mild assumption for complex, high-dimensional models. While this analysis provides an approximation rather than a strict bound, it formalizes the core intuition that reward variance acts as a natural, implicit regularizer, offering a strong theoretical motivation for the empirical stability of RFT in continual post-training. ∎
Appendix C Pseudo Code of RIF-RFT
Input: New task training set: $\mathcal{D}_{k}=\{(x_{i},y_{i}^{*})\}_{i=1}^{M}$; Current model policy: $\pi_{\theta}$; Number of rollouts per input: NN; Reward threshold: $\tau$
for each sample $(x_{i},y_{i}^{*})\in\mathcal{D}_{k}$ do
   Initialize $R_{\max}\leftarrow$-$\infty$;
 for $j=1$ to NN do
      Sample a response: $y_{ij}\sim\pi_{\theta}(\cdot\mid x_{i})$;
      Compute reward: $R(y_{ij})$;
      Update: $R_{\max}\leftarrow\max(R_{\max},R(y_{ij}))$;
 if $R_{\max}>\tau$ then
      Add $(x_{i},y_{i}^{*})$ to $\mathcal{D}_{k}^{\text{filt}}$;
Output: Updated model $\pi_{\theta^{\prime}}$
Algorithm 1 Rollout-based Instance Filtering for RFT (RIF-RFT)
Appendix D Statement on the Use of LLM
We disclose that LLMs were used to assist with writing. Their role was limited to improving the grammar and overall readability of the manuscript. The core research ideas, experimental results, and scientific claims presented are the work of the authors.